# CLAP: Isolating Content from Style through Contrastive Learning with Augmented Prompts

### v5

@@ B.1 Synthetic Image Generation
We utilize the Stable Diffusion v2.1 model [39] for generating synthetic image data used in our comparative experiments. Specifically, we adopt the implementation provided via the Stable Diffusion v2-1 Model Card on Hugging Face [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1. For each class across the four benchmark datasets, a total of 480 images are synthesized by applying our structured template prompts as input to the diffusion model. All generated images are rendered at a resolution of $512\times 512$ pixels. Representative examples of the synthetic images, along with their corresponding text prompts, are presented in fig. 7.
@@
Figure 7: Representative examples of synthetic images generated using Stable Diffusion v2.1, paired with their associated template-based prompts.
@@
Table 11: Zero-shot performance on VLCS across varied augmentation combinations and prompt sources. We evaluate average top-1 accuracy (%) for each domain using CLIP and CLAP variants trained with different prompt augmentations: ① Object Size Deletion, ② Object Color Deletion, ③ Image Type Deletion, ④ Art Style Deletion, ⑤ Prompt Order Swapping, and ⑥ Gaussian Noise Addition. Rightmost columns show EDA-based augmentations: LLM, Random, PromptStyler (Pr.St.), and Template.
Method Domain Avg. top-1 accuracy (%) (↑) on VLCS
CLIP (base) ①②③ ①②③ ①②③ ①②③ ③④⑤ ①②⑤ LLM Random Pr.St. Template
④⑤⑥ ④⑤ ④⑥
ZS(C) C 99.7 99.9 99.8 99.9 99.8 99.9 99.9 97.9 99.7 99.9 99.9
L 61.8 66.6 62.3 67.0 62.2 66.2 67.7 66.2 69.0 67.3 66.5
S 70.1 78.1 75.5 78.0 74.3 78.5 78.0 73.2 76.9 73.5 76.9
V 73.9 82.8 80.6 83.2 79.3 82.7 84.9 72.6 81.8 81.8 81.9
ZS(CP) C 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.8 99.9 99.9 99.9
L 69.9 69.3 67.9 69.6 68.4 70.0 70.4 69.3 70.4 71.2 69.7
S 73.3 77.6 76.4 76.7 75.9 78.8 77.2 76.2 75.2 75.1 78.0
V 84.8 85.3 84.0 85.3 84.2 85.1 86.0 77.0 84.2 86.0 84.6
ZS(PC) C 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9
L 70.2 70.0 68.0 70.1 68.5 70.0 70.7 67.5 70.6 71.8 70.0
S 73.6 76.6 75.6 76.0 74.8 77.8 76.9 76.9 75.1 74.9 78.2
V 86.1 85.7 84.7 85.7 84.5 85.5 86.2 78.2 84.6 86.8 84.8
ZS(NC) C 87.0 99.8 99.6 99.8 99.4 99.7 99.9 95.3 98.6 99.6 99.8
L 55.9 65.2 61.3 65.6 60.5 65.4 65.9 63.0 66.7 64.0 64.7
S 61.4 75.6 70.3 75.2 68.3 74.9 75.3 68.9 73.3 69.8 73.0
V 68.9 80.1 75.2 80.4 73.8 79.4 82.9 69.3 79.6 77.2 78.6
Table 12: Ablation study on hyperparameters for CLAP using ViT-B/16 on the VLCS dataset. We report domain-wise average top-1 accuracy (%) under three prompt settings—ZS(C), ZS(CP), and ZS(PC)—by varying: (1) $\tau$ (temperature), (2) latent dimension size, and (3) $\alpha$ (residual weight).
Hyper- Value Average Top-1 Accuracy (%) on VLCS (↑)
@@ C L S V C L S V C L S V
B.2 LLM Prompt Generation
We employ ChatGPT-3.5 [3] to generate the LLM-based prompts used in our comparative analysis of prompt sources. The procedure, illustrated in the following toolbox, involves prompting ChatGPT-3.5 to produce semantically rich, stylistically varied textual descriptions for specific class names. For each class, we generate 120 distinct samples. table 13 shows representative examples of the generated prompts.
@@ B.2 LLM Prompt Generation
B.3 Experiments on ViT-L/14
We refine the output dimension to match the ViT-L/14 input dimension of 768. The latent dimensions are set to 448 for PACS, 640 for VLCS, and 768 for both OfficeHome and DomainNet. The inference weighting parameter $\alpha$ is configured as 0.1 for PACS, 0.03 for VLCS, 0.14 for OfficeHome, and 0.2 for DomainNet. All other training configurations remain consistent with those used in the ViT-B/16 experiments. The training setup for Im.Aug mirrors that of CLAP for each dataset, with $\alpha$ set to 0.1 for PACS and 0.03 for the other three datasets.
Table 14 presents the zero-shot results for the ViT-L/14 model across four distinct prompts, following the same evaluation protocol as in the ViT-B/16 experiments. These results demonstrate that CLAP is more effective than Im.Aug in enhancing zero-shot performance. Furthermore, table 15 shows that CLAP substantially reduces performance variance across prompts, further validating its improvements over CLIP across multiple model scales. Detailed domain-level results are reported in table 16, providing a comprehensive breakdown of performance by domain.
B.4 Experiments on ResNet50x16
Table 17 showcases the zero-shot results for ResNet50x16 model across different prompts, substantiating that CLAP is more effective than Im.Aug in refining CLIP features. Moreover, table 18 illustrates that both Im.Aug and CLAP reduce variations in zero-shot performance across different prompts, with the improvement of CLAP being more significant. The results validate our approach across different model scales, including both ViT-based and CNN-based structures. Domain-level results are detailed in table 19.
Appendix C Discussion
C.1 Rationale behind CLAP’s Foundation on CLIP
C.2 Impact of Image and Text Augmentations
@@
Table 15: Prompt robustness analysis with CLIP pre-trained ViT-L/14 model. CLAP reduces the variance in zero-shot performance across different prompts.
@@
Table 16: Domain-level zero-shot results for the ViT-L/14 model across test datasets. CLAP consistently improves performance across all domains and prompt types, demonstrating its robustness and generalization capabilities.
@@
Table 17: Zero-shot performance with the CLIP pre-trained ResNet50x16 model. CLAP consistently enhances performance across all datasets, validating its effectiveness.
@@
Table 18: CLAP consistently reduces variance in zero-shot performance across different prompts using the CLIP pre-trained ResNet50x16 model, further validating its robustness and effectiveness.
@@
Table 19: Domain-level zero-shot performance using ResNet50x16 across test datasets. CLAP consistently improves upon CLIP and Im.Aug across domains and prompt types (ZS(C), ZS(CP), ZS(PC), ZS(NC)), demonstrating its generalizability and robustness.

### v6

@@ 0.B.1 Synthetic Image Generation
We employ the stable diffusion [39] v2.1 model for generating synthetic images used in our comparing experiments, specifically utilizing the Stable Diffusion v2-1 Model Card available on Hugging Face [^3] https://huggingface.co/stabilityai/stable-diffusion-2-1. For each class across the four datasets, we produce 480 images using our synthetic template prompts as input for the stable diffusion model. All generated images are of $512\times 512$ resolution. Examples of these synthetic images alongside their corresponding text prompts are displayed in Fig. 7.
@@
Figure 7: Examples of synthetic images created with SDv2.1 and associated prompts.
@@
Table 11: Zero-Shot Performance on VLCS Dataset Across Varied Augmentation Combinations and Prompt Sources: ① Random Object Size Deletion, ② Random Object Color Deletion, ③ Random Image Type Deletion, ④ Random Art Style Deletion, ⑤ Random Swapping Order, ⑥ Addition of Gaussian Noise.
Method Domains Avg. top-1 acc. (%) (↑) of different augmentations and prompts on VLCS
CLIP (base) ①②③ ①②③ ①②③ ①②③ ③④⑤ ①②⑤ EDA
④⑤⑥ ④⑤ ④⑥  LLM Rand. Pr.St. Temp.
ZS(C) C. 99.7 99.9 99.8 99.9 99.8 99.9 99.9 97.9 99.7 99.9 99.9
L. 61.8 66.6 62.3 67.0 62.2 66.2 67.7 66.2 69.0 67.3 66.5
S. 70.1 78.1 75.5 78.0 74.3 78.5 78.0 73.2 76.9 73.5 76.9
V. 73.9 82.8 80.6 83.2 79.3 82.7 84.9 72.6 81.8 81.8 81.9
ZS(CP) C. 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.8 99.9 99.9 99.9
L. 69.9 69.3 67.9 69.6 68.4 70.0 70.4 69.3 70.4 71.2 69.7
S. 73.3 77.6 76.4 76.7 75.9 78.8 77.2 76.2 75.2 75.1 78.0
V. 84.8 85.3 84.0 85.3 84.2 85.1 86.0 77.0 84.2 86.0 84.6
ZS(PC) C. 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9
L. 70.2 70.0 68.0 70.1 68.5 70.0 70.7 67.5 70.6 71.8 70.0
S. 73.6 76.6 75.6 76.0 74.8 77.8 76.9 76.9 75.1 74.9 78.2
V. 86.1 85.7 84.7 85.7 84.5 85.5 86.2 78.2 84.6 86.8 84.8
ZS(NC) C. 87.0 99.8 99.6 99.8 99.4 99.7 99.9 95.3 98.6 99.6 99.8
L. 55.9 65.2 61.3 65.6 60.5 65.4 65.9 63.0 66.7 64.0 64.7
S. 61.4 75.6 70.3 75.2 68.3 74.9 75.3 68.9 73.3 69.8 73.0
V. 68.9 80.1 75.2 80.4 73.8 79.4 82.9 69.3 79.6 77.2 78.6
Table 12: Ablative study of hyperparameters on VLCS dataset using ViT-B/16 model.
Hyper- Value Avg. top-1 acc. (%) (↑) using ViT-B/16 on VLCS dataset
@@
0.B.2 LLM Prompts Generation
We employ ChatGPT-3.5 [3] to generate the LLM-based prompts used in our comparative analysis of prompt sources. The procedure, illustrated in the following toolbox, involves prompting ChatGPT-3.5 to produce semantically rich, stylistically varied textual descriptions for specific class names. For each class, we generate 120 distinct samples. Tab. 13 shows representative examples of the generated prompts.
@@ 0.B.2 LLM Prompts Generation
Appendix 0.C Experiments on Other CLIP Model Scales
0.C.1 Experiments on ViT-L/14
We refined the output dimension to align with the input dimension of 768. The chosen latent dimensions were 448 and 640 for PACS and VLCS, respectively, and 768 for both OfficeHome and DomainNet. The inference weighting $\alpha$ was set to 0.1 for PACS, 0.03 for VLCS, 0.14 for OfficeHome, and 0.2 for DomainNet. All other training configurations remained consistent with the ViT-B/16 experiments across each dataset. The training configuration for Im.Aug was set the same as CLAP for each dataset, with the inference weighting $\alpha$ being 0.1 for PACS and 0.03 for the other three datasets.
Table 14 showcases the zero-shot results for the ViT-L/14 model using four distinct prompts, following the protocol established for the ViT-B/16 experiments. These results demonstrate that CLAP is more efficient than Im.Aug in enhancing zero-shot performance. Moreover, Tab. 15 illustrates that CLAP significantly reduces variations in zero-shot performance across different prompts, thereby confirming CLAP’s performance improvements over CLIP across a range of model sizes. Detailed domain-level results are presented in Tab. 16, offering an in-depth analysis.
@@ 0.C.1 Experiments on ViT-L/14
Table 15: CLAP reduces the variance in zero-shot performance across different prompts with CLIP pre-trained ViT-L/14 model.
@@
Table 16: Domain-level zero-shot results of the ViT-L/14 model on the test datasets.
@@
0.C.2 Experiments on ResNet50x16
Table 17 showcases the zero-shot results for ResNet50x16 model across different prompts, substantiating that CLAP is more effective than Im.Aug in refining CLIP features. Moreover, Tab. 18 illustrates that both Im.Aug and CLAP reduce variations in zero-shot performance across different prompts, with the improvement of CLAP being more significant. The results validate our approach across different model scales, including both ViT-based and CNN-based structures. Domain-level results are detailed in Tab. 19.
Table 17: Zero-shot performance with CLIP pre-trained ResNet50x16 model. CLAP demonstrates consistent enhancement across all datasets, validating its effectiveness.
@@ 0.C.2 Experiments on ResNet50x16
Table 18: CLAP consistently reduces variances in zero-shot performance across different prompts with CLIP pre-trained ResNet50x16 model, validating its effectiveness.
@@
Table 19: Domain-level zero-shot results using RestNet50x16 on the test datasets.
@@
Appendix 0.D Discussion
0.D.1 Rationale behind CLAP’s Foundation on CLIP
0.D.2 Impact of Image and Text Augmentations